OpenStudy (anonymous):

Given the arithmetic sequence 3, 7, 11, 15, …, what is a1, the first term? What is d, the common difference? Find the 151st term of this sequence. What formula did you use? Be sure to show your work.

OpenStudy (loser66):

Why did you post it twice? @peachpi is helping you.

OpenStudy (anonymous):

A1 is 3. D value is7-3=4 To find the 151st value you use An=a1+(n-p)d A151=3+(151-1)x4 A151=3+150x4 A151=3+600 A151=603